Listed here are the total costs associated with the production of 1,000 drum sets manufactured by TrueBeat. The drum sets sell for $540 each.   

Costs

Plastic for casing—$15,000

Wages of assembly workers—$90,000

Property taxes on factory—$8,000

Accounting staff salaries—$42,000

Drum stands (1,000 stands purchased)—$38,000

Rent cost of equipment for sales staff—$36,000

Upper management salaries—$205,000

Annual flat fee for factory maintenance service—$20,000

Sales commissions—$25 per unit

Machinery depreciation, straight-line—$35,000

  1. Increase Fixed Costs: TrueBeat’s management is also considering purchasing additional advertising for $40,000. Marketing studies suggest that sales will increase by 200 units a year. Should the company purchase the additional advertising? Your answer should use complete sentences and contain 30 to 50 words. Support your answer with calculations of at least 2 of the following items. (a) Net Operating Income, (b) Break-even in units or dollars,(c) target sales in units or dollars to maintain the same Net Income as originally assumed for year 4. You may also use other calculations that you feel are relevant.
Answer #1
Let us find the contribution margin , Net Income
and BEP units at 1000 units level:
Variable costing Income Statement Per unit $ Total Amount $
Units Sold                   1,000
Sales Revenue $              540 $          540,000
Variable costs
Plastic for casing $                15 $            15,000
Wages for Assembly workers $                90 $            90,000
Drum Stands $                38 $            38,000
Sales Commission $                25 $            25,000
Total Variable cost $              168 $          168,000
Contribution Margin $              372 $          372,000
Contribution Margin % 68.89% 68.89%
Fixed Costs
Property Tax of Factory $              8,000
Accounting staff salaries $            42,000
Rent for sales Equipment $            36,000
Machine depreciation $            35,000
Annual Maintenance fees -flat $            20,000
Management Salaries $          205,000
Total Fixed costs $          346,000
Net Operating Income /(Loss) $            26,000
BEP Analysis
Contribution Margin 68.89%
Total Fixed costs $      346,000
BEP Sales =346000/68.89%= $      502,258
Units Sales Price = $              540
BEP Units =502258/540=               930.1
Deciding on Additional Advertisisng cost
Incremental Sales per year in units                  200
Contribution Margin per unit= $              372
Incremental Contribution = $         74,400
Incremental Fixed cost of Advertsing $         40,000
Incremental Operating Income $         34,400
The company should opt for increasing the
advertsing cost by $40,000 annually.
The incremental Net Income that will be generated
by the incremental contribution from the 200 units
of incremental sales is $34,400 more than the
additional advertising expense planned .
So the company should purchase the additional
advertising.